OpenStudy (anonymous):

Lim x to infinity x^(1/2)*sin(((x+2)^1/2) - ((x-2)^1/2))

OpenStudy (anonymous):

is that:\[\sqrt{x}*\sin(\sqrt{x+2}-\sqrt{x-2})\]?

OpenStudy (anonymous):

\[\lim_{x \rightarrow \infty} \sqrt{x} \sin (\sqrt{x+2}-\sqrt{x-2})\]

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

first simplify the term under sin.rationalise it

OpenStudy (anonymous):

oops.sorry,that would complicate it further.no need to do that

OpenStudy (anonymous):

why? thats a good point for start..

OpenStudy (anonymous):

after doing that we just need to apply\[\sin z \approx z \]when z approaches to zero

OpenStudy (anonymous):

if im not wrong :)

OpenStudy (anonymous):

well the answer is 2 so idk

OpenStudy (anonymous):

i just dont know how to get there

OpenStudy (anonymous):

thats right answer is 2

OpenStudy (anonymous):

ok see starting with the hint MotherofGod gave u\[\lim_{x \rightarrow \infty} \sqrt{x} \sin (\sqrt{x+2}-\sqrt{x-2})=\lim_{x \rightarrow \infty} \sqrt{x} \sin (\frac{4}{\sqrt{x+2}+\sqrt{x-2}})\]because\[\sqrt{x+2}-\sqrt{x-2}=\sqrt{x+2}-\sqrt{x-2}*\frac{\sqrt{x+2}+\sqrt{x-2}}{\sqrt{x+2}+\sqrt{x-2}}=\frac{4}{\sqrt{x+2}+\sqrt{x-2}}\]

OpenStudy (anonymous):

\[\frac{4}{\sqrt{x+2}+\sqrt{x-2}}\]goes to zero as x goes to infty so\[\sin(\frac{4}{\sqrt{x+2}+\sqrt{x-2}}) \approx \frac{4}{\sqrt{x+2}+\sqrt{x-2}}\]rewrite the limit\[\lim_{x \rightarrow \infty} \sqrt{x} \sin (\sqrt{x+2}-\sqrt{x-2})=\lim_{x \rightarrow \infty}\frac{4\sqrt{x}} {\sqrt{x+2}+\sqrt{x-2}}=\frac{4\sqrt{x}}{2\sqrt{x}}=2\]

OpenStudy (anonymous):

(x+2)^1/2 + (x-2)^2 = 2x^1/2 ???

OpenStudy (anonymous):

as x goes to infty 2 and -2 will be neglegible in comparison with x

OpenStudy (anonymous):

yeahh!!! Great!!! Thx a lot

OpenStudy (anonymous):

very welcome :)